What is the magnitude of the electric field at a point midway between a −5.3μC and a +5.4μC charge 8.8cm apart? Assume no other charges are nearby.
Answer #1

Here , distance of midpoint from the charges

d = 8.8/2 cm = 4.4 cm

d = 0.044 m

Now, for the electric field

magnitude of electric field = k * (q1 + q2)/d^2

magnitude of electric field = 9 *10^9 * (5.3 + 5.4) *10^-6/0.044^2

calculating

magnitude of electric field = 4.97 *107 N/C

the magnitude of electric field is 4.97 *107 N/C